OpenStudy (anonymous):

Factor z3 - 2z2 + 9z - 18. (z2 + 9)(z - 2) (z2 + 2)(z - 9) (z2 - 9)(z - 2)

OpenStudy (anonymous):

1st one

OpenStudy (anonymous):

use your common sense.

OpenStudy (anonymous):

\[(z-2) \left(z^2+9\right)=z^3-2 z^2+9 z-18 \]